OpenStudy (anonymous):

Tena is five years older than her sister Tammy. The product of their ages is 234. How old is Tammy?

OpenStudy (anonymous):

a..12 b. 10 c. 13 d. 24

OpenStudy (anonymous):

Tammy's age is \(x\), and Tena's age is \(x+5\). We know that \(x\cdot(x+5) = 234\).

OpenStudy (anonymous):

This will simplify to: \[ x^2+5x-234 = 0 \]You can solve this like any other quadratic equation.

OpenStudy (anonymous):

Still figuring it out. i almost got it.

OpenStudy (anonymous):

Is Tammy 13? Because I took 13 * 18 and that equaled 234.

OpenStudy (anonymous):

If that worked, then yes.

OpenStudy (anonymous):

Did you guess, or did you solve it with algebra?

OpenStudy (anonymous):

I solved it by taking the four answers I had and added the 5 years to each and then just multiplied them.

OpenStudy (anonymous):

Thank u for your help